OpenStudy (anonymous):

Describe in detail how you would construct a 95% confidence interval for a set of 30 data points whose mean is 20 and population standard deviation is 3.

OpenStudy (displayerror):

n = 30, so we can say we have a large enough sample to use the normal distribution (z-table) Write out the given values: sample mean, \[\bar{X} = 20\] and sample standard deviation, \[\sigma = 3\] Recall that the general expression for a confidence interval with large enough sample and a known population standard deviation is given by \[\bar{X} \pm Z_{\frac{\alpha}{2}} \frac{\sigma}{\sqrt{n}}\] Now plug in your given values, where \[\alpha\] is 0.05 (5%). Find the value on the table, plug it in, and you're done!

OpenStudy (anonymous):

thnx man yo helped alot